OpenStudy (anonymous):

If you deposit $1000 in a savings account with an interest rate of r compounded annually, then the balance in the account after 3 years is given by the function B(c) = 1000(1 + r)3, where r is written as a decimal. What interest rate will yield a balance of $1100 after 3 years? a. 3.23% c. 3.63% b. 3.33% d. 3.66%

OpenStudy (anonymous):

help me out please!

OpenStudy (anonymous):

The amount in the savings account is given by the function. So, we just set $1100 equal to the equation and solve for r: \[1100=1000(1+r)^3\]Divide both sides by 1000: \[1.1=(1+r)^3\]Cube root both sides: \[1.03228=1+r\]Subtract r from both sides: \[r=.03228\]

OpenStudy (anonymous):

thanks!